# Revisiting the Orbital Dynamics of the Hot Jupiter WASP–12b with New Transit Times

### v1

@@ 5.1 Goodness-of-fit metrics
$\Delta\mathrm{BIC}=\mathrm{BIC}_{\mathrm{lin}}-\mathrm{BIC}_{\mathrm{quad}}=1066.12.$ (6)
@@
$B=\exp\left(\frac{\Delta\mathrm{BIC}}{2}\right)\approx 3.20\times 10^{231}.$ (7)
@@
$\Delta\mathrm{AIC}=\mathrm{AIC}_{\mathrm{lin}}-\mathrm{AIC}_{\mathrm{quad}}=1070.08,$ (8)
which further supports orbital decay as the most plausible explanation of the observed transit-timing variations. These findings are consistent with previous studies (G. Maciejewski et al. 2016; K. C. Patra et al. 2017; S. W. Yee et al. 2020).
6 Discussions
6.1 Comparison of the orbital decay rate with previous studies
Earlier transit-timing studies (K. C. Patra et al. 2017; G. Maciejewski et al. 2018; J. D. Turner et al. 2021) reported compelling evidence for a decreasing orbital period. By incorporating a substantial set of newly available transit timings from ETD and ExoClock, along with additional ground-based photometric observations that significantly extend the temporal baseline, we performed an updated ephemeris analysis to obtain a more precise estimate of the orbital decay rate.
@@ 6.1 Comparison of the orbital decay rate with previous studies
$\dot{P}=-30.31\pm 0.92~\mathrm{ms\,yr^{-1}},$ (9)
which is fully consistent with recent measurements and lies within the $1\sigma$ confidence interval of several reported values. For instance, N. Sodickson & S. Grunblatt (2025), who included the recent TESS sectors together with the IW22 data, reported $\dot{P}=-30.85\pm 0.82~\mathrm{ms\,yr^{-1}}$. Similarly, P. Leonardi et al. (2024), incorporating both spectroscopic and photometric data, obtained $\dot{P}=-30.72\pm 2.67~\mathrm{ms\,yr^{-1}}$. Using only TESS data,E. S. Ivshina & J. N. Winn (2022) reported $\dot{P}=-30.27\pm 1.11~\mathrm{ms\,yr^{-1}}$.
Our derived value is also in agreement with earlier results, lying within $\sim 1.15\sigma$ of the estimate $\dot{P}=-2.56\pm 4.0~\mathrm{ms\,yr^{-1}}$ reported by G. Maciejewski et al. (2016). Furthermore, it remains consistent with the results of K. C. Patra et al. (2017) and S. W. Yee et al. (2020), who included occultation data along with transit data in their analyses.
@@
This work -30.31 $\pm$ 0.92
N. Sodickson & S. Grunblatt (2025) -30.85 $\pm$ 0.82
Z. Shen (2024) -26.31 $\pm$ 0.90
P. Leonardi et al. (2024) -30.72 $\pm$ 2.67
E. R. Adams et al. (2024) -29.8 $\pm$ 1.6
E. Alvarado et al. (2024) -29.5 $\pm$ 1.0
E. S. Ivshina & J. N. Winn (2022) -30.27 $\pm$ 1.11
I. Wong et al. (2022) -29.81 $\pm$ 0.94
L. Bai et al. (2022) -37.14 $\pm$ 1.31
J. D. Turner et al. (2021) -32.53 $\pm$ 1.62
S. W. Yee et al. (2020) -29 $\pm$ 2
K. C. Patra et al. (2017) -29 $\pm$ 3
G. Maciejewski et al. (2016) -25.6 $\pm$ 4.0
6.2 Calculation of Orbital Decay Timescale of WASP-12b
The orbital decay timescale ($T_{d}$) quantifies the characteristic time over which a hot Jupiter gradually spirals inward due to tidal interactions with its host star, ultimately leading to its engulfment. This timescale provides an important measure of the long-term dynamical evolution of close-in giant planets. Using the orbital period ($P$) and the orbital decay rate ($\dot{P}$) derived in Section 5, we estimate the decay timescale as
$T_{d}=\frac{P}{\dot{P}}\approx 3.1\ \text{Myr}.$ (10)
Our derived value of $T_{d}$ is in good agreement with the result reported by S. W. Yee et al. (2020), who obtained $T_{d}=3.25$ Myr, and is slightly higher than the estimate of $T_{d}=2.90$ Myr derived by J. D. Turner et al. (2021). This close correspondence reinforces the reliability of the measured orbital decay rate and suggests a consistently rapid inward migration for WASP-12b over astrophysical timescales.
6.3 Calculation of Stellar Tidal Quality Factor
The identification of close-in hot Jupiters has renewed focus on the role of tidal interactions in governing stellar energy dissipation (R. I. Dawson & J. A. Johnson 2018). Central to this process is the tidal quality factor, ${Q}{\ast}$, which serves as a measure of how effectively a star dissipates tidal energy. Formally, ${Q}{\ast}$ is expressed as the ratio between the maximum energy stored in the tidal distortion of the star during an orbital cycle and the total energy lost to frictional dissipation in the same interval (see, e.g., Eq. (2.19) of J. P. Zahn 2008). This quantity is fundamental in determining the characteristic timescales over which star-planet tidal interactions affect the stellar spin evolution and the orbital dynamics of the companion.
Since WASP-12b exhibits a markedly negative orbital decay rate (see Section 5), the reduction in its orbital period is most plausibly explained by tidal dissipation within the host star. On this basis, we determine the modified stellar tidal quality factor ($Q^{\prime}_{\ast}$), a dimensionless parameter that empirically characterizes the efficiency of tidal kinetic energy dissipation within the star. To compute $Q^{\prime}_{\star}$ for the WASP-12 system, we adopt the formalism of the modified constant phase-lag model (P. Goldreich & S. Soter 1966), following the methodology of K. C. Patra et al. (2017), A. N. Wilkins et al. (2017), and G. Maciejewski et al. (2018). The relevant expression (Equation 7) is given by
@@ 6.3 Calculation of Stellar Tidal Quality Factor
Adopting $M_{p}/M_{\star}$ and $a/R_{\star}=3.061$ from I. Wong et al. (2022), and substituting the measured $\dot{P}$ (from Section 4.3), we infer $Q^{\prime}_{\star}\approx 1.61\times 10^{5}$ for WASP-12. This value lies within the range previously inferred for hot-Jupiter host stars ($10^{5}$$10^{6.5}$; B. Jackson et al. 2008; N. Husnoo et al. 2012; A. J. Barker 2020), for binary systems ($10^{5}$$10^{7}$; S. Meibom & R. D. Mathieu 2005; G. I. Ogilvie & D. N. C. Lin 2007; A. F. Lanza 2010; S. Meibom et al. 2015), and for transiting giant planets ($10^{4}$$10^{8}$; A. S. Bonomo et al. 2017). This value is also consistent with estimates for stars hosting gas-giant planets on ultra-short orbital periods, which typically exhibit $Q^{\prime}_{\star}$ values in the range of $10^{5}$$10^{7}$ (K. Penev et al. 2018). Furthermore, the modified stellar tidal quality factor ($Q^{\prime}_{\star}$) obtained in our analysis is of the same order of magnitude as $Q^{\prime}_{\star}=4.3\times 10^{5}$, as derived from the theoretical models of R. Essick & N. N. Weinberg (2016) for solar-type host stars.
While our derived $Q^{\prime}_{\star}$ agrees with these results, it is 1–2 orders of magnitude smaller than the typical values reported for Sun-like primaries in eclipsing binaries ($\sim 10^{7.8}$; R. Patel & K. Penev 2022), simplified tidal evolution models ($10^{7.5}$$10^{8.5}$; B. M. S. Hansen 2010), and hot Jupiters in dynamical and equilibrium tide regimes ($10^{7.3}$$10^{8.3}$; A. Collier Cameron & M. Jardine 2018). Our estimate of $Q^{\prime}_{\star}=1.61\times 10^{5}$ therefore implies efficient tidal dissipation and rapid orbital decay. Moreover, this result falls within the $1\sigma$ lower confidence bound of the value reported by N. Sodickson & S. Grunblatt (2025), namely $Q^{\prime}_{\star}=1.64\times 10^{5}$. A comparison of our result with values from earlier studies is summarized in Table 12.
@@
$1.61\times 10^{5}$ This Work
$1.64\times 10^{5}$ N. Sodickson & S. Grunblatt (2025)
$2.13\times 10^{5}$ P. Leonardi et al. (2024)
$1.7\times 10^{5}$ L.-C. Yeh et al. (2024)
$1.6\times 10^{5}$ E. Alvarado et al. (2024)
$1.50\times 10^{5}$ I. Wong et al. (2022)
$1.39\times 10^{5}$ J. D. Turner et al. (2021)
$1.75\times 10^{5}$ S. W. Yee et al. (2020)
$2.5\times 10^{5}$ G. Maciejewski et al. (2016)
$4.3\times 10^{5}$ R. Essick & N. N. Weinberg (2016)
6.4 Calculation of the shift in transit time and the Remaining Lifetime
The expected shift in the transit arrival time of an exoplanet, denoted as $T_{\mathrm{shift}}$, corresponds to the predicted variation in transit timing as observed from Earth, arising from perturbations in the planet’s orbital dynamics. To estimate the anticipated timing shift for WASP-19b under the influence of orbital decay, we employed Equation (7) of J. L. Birkby et al. (2014):
@@ 6.4 Calculation of the shift in transit time and the Remaining Lifetime
where $dn/dT$ represents the present rate of change in the orbital frequency of the planet. The orbital period ($P$) was adopted for the orbital decay model from the Table 10 . Using the calculated value $dn/dT\approx 6.79\times 10^{-19}~\mathrm{rad\,s^{-2}}$, corresponding to a modified stellar tidal quality factor of $Q^{\prime}_{\star}=1.6\times 10^{5}$, the predicted transit timing shift after $T=17$ yr is found to be $T_{\mathrm{shift}}\approx 1465.96$ s. This prediction can be tested and refined through future follow-up transit monitoring.
The remaining lifetime of a hot Jupiter corresponds to the timescale over which its orbit decays sufficiently for the planet to spiral inward and eventually merge with its host star. This orbital decay is predominantly governed by stellar tidal dissipation, commonly characterized by the modified stellar tidal quality factor ($Q^{\prime}_{\star}$), which quantifies the efficiency of tidal energy dissipation in the star. Consequently, the orbital evolution timescale of hot Jupiters is strongly dependent on the adopted value of $Q^{\prime}_{\star}$. By substituting the appropriate $Q^{\prime}_{\star}$ value (see Section 6.1) along with other system parameters from I. Wong et al. (2022) into Equation (5) of B. Levrard et al. (2009):
@@
we estimate the remaining lifetime of WASP-12b to be $\sim 0.44$ Myr, where $n=2\pi/P$ denotes the mean orbital motion of the planet.
6.5 Estimation of Planetary Love Number
From the measured orbital eccentricity of WASP-12b ($e=0.0031^{+0.0000327}_{-0.0000478}$; see Section 5), one possible explanation for the observed TTVs is apsidal precession. However, tidal evolution theory predicts that hot Jupiters should circularize on timescales much shorter than the ages of their host stars (B. Levrard et al. 2007; R. I. Dawson & J. A. Johnson 2018). Assuming a planetary tidal dissipation factor of $Q_{p}\sim 10^{6}$ and applying the formalism of K. C. Patra et al. (2017), we estimate the tidal circularization timescale of WASP-12b to be $\sim$ 0.34  Myr. This timescale is several orders of magnitude shorter than the estimated stellar age (WASP-12: $\sim$3.05  Gyr, P. Leonardi et al. 2024), indicating that the planet’s orbit should have long been circularized. Any residual orbital eccentricity would therefore require additional mechanisms, such as ongoing perturbations or a reevaluation of tidal dissipation efficiency, to be sustained (L. G. Bouma et al. 2019; G. Maciejewski et al. 2021).
A plausible explanation is that apsidal precession is influenced by the planet’s internal structure, as proposed by D. Ragozzine & A. S. Wolf (2009), since the precession rate depends on the tidal Love number ($k_{p}$), which encodes information about the planet’s internal density distribution. Following the formulation of K. C. Patra et al. (2017):
@@ 6.5 Estimation of Planetary Love Number
we estimated $k_{p}$ by substituting the values of $d\omega/dE$ (Table 8) together with other system parameters from I. Wong et al. (2022). This yields $k_{p}=0.66\pm 0.28$, a value consistent with Jupiter’s tidal Love number ($k_{p}=0.59$; S. M. Wahl et al. 2016). This similarity suggests that the interior density profile of WASP-12b may be comparable to that of Jupiter, and hence apsidal precession cannot be fully excluded as another contributing factor to the observed TTVs. To test this hypothesis, continued monitoring of future transits and occultations will be essential.
7 Conclusions
@@
Previous studies of WASP-12b (G. Maciejewski et al., 2016; K. C. Patra et al., 2017; S. W. Yee et al., 2020) have revealed intriguing aspects of the system’s orbital dynamics. Early investigations reported the presence of short-term transit timing variations (TTVs) potentially caused by an additional body (G. Maciejewski et al., 2011), while later studies identified evidence of long-term TTVs. Specifically, G. Maciejewski et al. (2016); K. C. Patra et al. (2017); G. Maciejewski et al. (2018) observed a declining trend in transit timings indicative of orbital decay, whereas A. Bailey & J. Goodman (2019) proposed apsidal precession as a plausible mechanism. More recently, S. W. Yee et al. (2020) provided strong evidence supporting orbital decay in the WASP-12b system. Detecting such orbital evolution requires long-term, high-precision monitoring. In this work, we combined newly acquired ground-based photometric data from 7 nights, obtained at various observatories worldwide, with publicly available light curves from the literature, ETD, ExoClock, and high-cadence, high-precision space-based observations from TESS to construct an extensive dataset for re-examining this hot Jupiter.
@@
We included a total of 119 complete transit light curves of WASP-12b observed by TESS across six sectors (20, 43, 44, 45, 71, and 72). To extend the observational time baseline, we incorporated an additional 250 high-quality transit light curves from public databases, including ETD and ExoClock, along with 108 publicly available complete transit observations. Furthermore, 26 transit light curves were obtained via private communication with the respective authors. In total, 391 light curves were modeled to refine the system’s physical and orbital parameters and to determine individual mid-transit times. To ensure uniformity and precision, all light curves were analyzed using a consistent methodology for deriving the mid-transit times.
@@
Our transit-timing analysis indicates that the orbit of WASP-12b is decaying at a rate of $30.31\,\mathrm{ms\,yr^{-1}}$, corresponding to a modified stellar tidal quality factor of $Q^{\prime}_{\star}=1.61\times 10^{5}$. Statistical diagnostics, including reduced $\chi^{2}$, BIC, and AIC, favor the orbital decay model as the most plausible explanation for the observed timing variations.
@@
Given the system’s small but non-negligible eccentricity ($e>0.003$), apsidal precession remains a viable alternative mechanism. To investigate this possibility and constrain the planet’s internal structure, we calculated the planetary tidal Love number, obtaining $k_{p}=0.66\pm 0.28$. This value is consistent with Jupiter’s Love number ($k_{p}=0.59$; S. M. Wahl et al. 2016), suggesting a similar internal density distribution for WASP-12b. Consequently, apsidal precession cannot be fully ruled out as another contributing factor, and continued high-precision monitoring of future transits with upcoming missions such as PLATO and Ariel, together with occultation measurements, will be crucial for distinguishing between orbital decay and apsidal precession as the dominant mechanism.
@@
This research is supported in part by the National Science and Technology Council (NSTC), Taiwan, under the Grants NSTC 113-2112-M-007-030, NSTC 114-2112-M-007-029, and NSTC 113-2115-M-007-008. This work utilizes new photometric observations obtained with the 0.61 m VASISTHA Telescope at the Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), Sitapur, Paschim Midnapore, operated by ICSP. We are grateful to Prof. Sandip Kumar Chakraborty for providing access to this facility and to the IERCOO staff, particularly Debasis Bhowmick and Mohit Singh Bisht, for their invaluable support during the observing runs. Additionally, this study incorporates observations from the 1.3 m Devasthal Fast Optical Telescope (DFOT) at the Devasthal Campus of ARIES, Nainital, and the 0.3 m AG Optical telescope at the Utah Desert Remote Observatory (UDRO) in Beryl, Utah. We thank the staff at DFOT and UDRO for their assistance. This work also makes use of data from the Transiting Exoplanet Survey Satellite (TESS) mission, publicly available through the Mikulski Archive for Space Telescopes (MAST). We gratefully acknowledge the TESS mission for its significant contribution to exoplanet science by providing high-precision transit photometric data. The specific observations analyzed in this study can be accessed via the TESS LCs–All Sectors page (MAST Team 2021 ), and funding for the mission is provided by NASA’s Explorer Program. Furthermore, this study utilized publicly available transit light curves from the Exoplanet Transit Database (ETD) and the ExoClock Project; we thank the contributors of these initiatives for making their data accessible. Currently, all the original lightcurves of ETD are available in VarAstro Server. We also acknowledge Hebb Lesbie, Efrain Alavarado, and Karen Collins for sharing their transit light curves through private communication, as well as all other authors who made their published light curves publicly available, which proved to be an invaluable resource for this study. Finally, this work made use of the NASA Exoplanet Archive, operated by the California Institute of Technology under contract with NASA as part of the Exoplanet Exploration Program.

### v2

@@ V.1 Goodness-of-fit metrics
$\Delta\mathrm{BIC}=\mathrm{BIC}_{\mathrm{lin}}-\mathrm{BIC}_{\mathrm{quad}}=1557.99.$ (7)
@@
$B=\exp\left(\frac{\Delta\mathrm{BIC}}{2}\right)\approx 1.58\times 10^{338}.$ (8)
@@
$\Delta\mathrm{AIC}=\mathrm{AIC}_{\mathrm{lin}}-\mathrm{AIC}_{\mathrm{quad}}=1165.98,$ (9)
which further supports orbital decay as the most plausible explanation of the observed transit-timing variations. These findings are consistent with previous studies (Maciejewski et al. 2016b; Patra et al. 2017; Yee et al. 2020).
VI Discussions
VI.1 Comparison of the orbital decay rate with previous studies
Earlier transit-timing studies (Patra et al. 2017; Maciejewski et al. 2018; Turner et al. 2021) reported compelling evidence for a decreasing orbital period. By incorporating a substantial set of newly available transit timings from ETD and ExoClock, along with additional ground-based photometric observations that significantly extend the temporal baseline, we performed an updated ephemeris analysis to obtain a more precise estimate of the orbital decay rate.
@@ VI.1 Comparison of the orbital decay rate with previous studies
$\dot{P}=-31.97\pm 0.80~\mathrm{ms\,yr^{-1}},$ (10)
which is fully consistent with recent measurements and lies within the $1\sigma$ confidence interval of several reported values. For instance, Sodickson and Grunblatt (2025), who included the recent TESS sectors together with the IW22 data, reported $\dot{P}=-30.85\pm 0.82~\mathrm{ms\,yr^{-1}}$. Similarly, Leonardi et al. (2024b), incorporating both spectroscopic and photometric data, obtained $\dot{P}=-30.72\pm 2.67~\mathrm{ms\,yr^{-1}}$. Using only TESS data,Ivshina and Winn (2022) reported $\dot{P}=-30.27\pm 1.11~\mathrm{ms\,yr^{-1}}$.
Our derived value is also in agreement with earlier results, lying within $\sim 1.4\sigma$ of the estimate $\dot{P}=-2.56\pm 4.0~\mathrm{ms\,yr^{-1}}$ reported by Maciejewski et al. (2016a). Furthermore, it remains consistent with the results of Patra et al. (2017) and Yee et al. (2020), who included occultation data along with transit data in their analyses.
@@
This work -31.97 $\pm$ 0.80
Sodickson and Grunblatt (2025) -30.85 $\pm$ 0.82
Shen (2024) -26.31 $\pm$ 0.90
Leonardi et al. (2024a) -30.72 $\pm$ 2.67
Adams et al. (2024) -29.8 $\pm$ 1.6
Alvarado et al. (2024) -29.5 $\pm$ 1.0
Ivshina and Winn (2022) -30.27 $\pm$ 1.11
Wong et al. (2022) -29.81 $\pm$ 0.94
Bai et al. (2022) -37.14 $\pm$ 1.31
Turner et al. (2021) -32.53 $\pm$ 1.62
Yee et al. (2020) -29 $\pm$ 2
Patra et al. (2017) -29 $\pm$ 3
Maciejewski et al. (2016b) -25.6 $\pm$ 4.0
VI.2 Calculation of Orbital Decay Timescale of WASP-12 b
The orbital decay timescale ($T_{d}$) quantifies the characteristic time over which a hot Jupiter gradually spirals inward due to tidal interactions with its host star, ultimately leading to its engulfment. This timescale provides an important measure of the long-term dynamical evolution of close-in giant planets. Using the orbital period ($P$) and the orbital decay rate ($\dot{P}$) derived in Section V, we estimate the decay timescale as
$T_{d}=\frac{P}{\dot{P}}\approx 2.94\ \text{Myr}.$ (11)
Our derived value of $T_{d}$ is fully consistent with $T_{d}=2.90$ Myr reported by Turner et al. (2021), and also with $T_{d}=3.25$ Myr obtained by Yee et al. (2020). This close correspondence reinforces the reliability of the measured orbital decay rate and suggests a consistently rapid inward migration for WASP-12 b over astrophysical timescales.
VI.3 Calculation of Stellar Tidal Quality Factor
The identification of close-in hot Jupiters has renewed focus on the role of tidal interactions in governing stellar energy dissipation (Dawson and Johnson 2018). Central to this process is the tidal quality factor, ${Q}{\ast}$, which serves as a measure of how effectively a star dissipates tidal energy. Formally, ${Q}{\ast}$ is expressed as the ratio between the maximum energy stored in the tidal distortion of the star during an orbital cycle and the total energy lost to frictional dissipation in the same interval (see, e.g., Eq. (2.19) of Zahn 2008). This quantity is fundamental in determining the characteristic timescales over which star-planet tidal interactions affect the stellar spin evolution and the orbital dynamics of the companion.
Since WASP-12b exhibits a markedly negative orbital decay rate (see Section V), the reduction in its orbital period is most plausibly explained by tidal dissipation within the host star. On this basis, we determine the modified stellar tidal quality factor ($Q^{\prime}_{\ast}$), a dimensionless parameter that empirically characterizes the efficiency of tidal kinetic energy dissipation within the star. To compute $Q^{\prime}_{\star}$ for the WASP-12 system, we adopt the formalism of the modified constant phase-lag model (Goldreich and Soter 1966), following the methodology of Patra et al. (2017), Wilkins et al. (2017), and Maciejewski et al. (2018). The relevant expression (Equation 7) is given by
@@ VI.3 Calculation of Stellar Tidal Quality Factor
Adopting $M_{p}/M_{\star}$ and $a/R_{\star}=3.061$ from Wong et al. (2022), and substituting the measured $\dot{P}$ (from Section V), we infer $Q^{\prime}_{\star}\approx(1.52\pm 0.038)\times 10^{5}$ for WASP-12. This value lies within the range previously inferred for hot-Jupiter host stars ($10^{5}$$10^{6.5}$; Jackson et al. 2008; Husnoo et al. 2012; Barker 2020), for binary systems ($10^{5}$$10^{7}$; Meibom and Mathieu 2005; Ogilvie and Lin 2007; Lanza 2010; Meibom et al. 2015), and for transiting giant planets ($10^{4}$$10^{8}$; Bonomo et al. 2017). This value is also consistent with estimates for stars hosting gas-giant planets on ultra-short orbital periods, which typically exhibit $Q^{\prime}_{\star}$ values in the range of $10^{5}$$10^{7}$ (Penev et al. 2018). Furthermore, the modified stellar tidal quality factor ($Q^{\prime}_{\star}$) obtained in our analysis is of the same order of magnitude as $Q^{\prime}_{\star}=4.3\times 10^{5}$, as derived from the theoretical models of Essick and Weinberg (2016) for solar-type host stars.
While our derived $Q^{\prime}_{\star}$ agrees with these results, it is 1–2 orders of magnitude smaller than the typical values reported for Sun-like primaries in eclipsing binaries ($\sim 10^{7.8}$; Patel and Penev 2022), simplified tidal evolution models ($10^{7.5}$$10^{8.5}$; Hansen 2010), and hot Jupiters in dynamical and equilibrium tide regimes ($10^{7.3}$$10^{8.3}$; Collier Cameron and Jardine 2018). Our estimate of $Q^{\prime}_{\star}=1.6\times 10^{5}$ therefore implies efficient tidal dissipation and rapid orbital decay. Moreover, this result falls within the $1\sigma$ lower confidence bound of the value reported by Sodickson and Grunblatt (2025), namely $Q^{\prime}_{\star}=1.64\times 10^{5}$. A comparison of our result with values from earlier studies is summarized in Table 11.
@@
$(1.52\pm 0.038)\times 10^{5}$ This Work
$(1.64\pm 0.04)\times 10^{5}$ Sodickson and Grunblatt (2025)
$(2.13\pm 0.18)\times 10^{5}$ Leonardi et al. (2024a)
$(~1.7\pm 0.0405\times 10^{5}$ Yeh et al. (2024)
$(1.7\pm 0.14)\times 10^{5}$ Akinsanmi et al. (2024)
$(1.6\pm 0.1)\times 10^{5}$ Alvarado et al. (2024)
$(1.5\pm 0.11)\times 10^{5}$ Wong et al. (2022)
$(1.39\pm 0.15)\times 10^{5}$ Turner et al. (2021)
$(1.75\pm 0.12)\times 10^{5}$ Yee et al. (2020)
$2.5\times 10^{5}$ Maciejewski et al. (2016b)a
$4.3\times 10^{5}$ Essick and Weinberg (2016)a
Notes: a It does not report the uncertainty (error bar) associated with the modified stellar tidal quality factor in its respective paper.
VI.4 Calculation of the shift in transit time and the Remaining Lifetime
The expected shift in the transit arrival time of an exoplanet, denoted as $T_{\mathrm{shift}}$, corresponds to the predicted variation in transit timing as observed from Earth, arising from perturbations in the planet’s orbital dynamics. To estimate the anticipated timing shift for WASP-12 b under the influence of orbital decay, we employed Equation (7) of Birkby et al. (2014):
@@ VI.4 Calculation of the shift in transit time and the Remaining Lifetime
where $dn/dT$ represents the present rate of change in the orbital frequency of the planet. The orbital period ($P$) was adopted for the orbital decay model from the Table 9 . Using the calculated value $dn/dT\approx 7.16\times 10^{-19}~\mathrm{rad\,s^{-2}}$, corresponding to a modified stellar tidal quality factor of $Q^{\prime}_{\star}=1.52\times 10^{5}$, the predicted transit timing shift after $T=17$ yr is found to be $T_{\mathrm{shift}}\approx 1546.02$ s. This prediction can be tested and refined through future follow-up transit monitoring.
The remaining lifetime of a hot Jupiter corresponds to the timescale over which its orbit decays sufficiently for the planet to spiral inward and eventually merge with its host star. This orbital decay is predominantly governed by stellar tidal dissipation, commonly characterized by the modified stellar tidal quality factor ($Q^{\prime}_{\star}$), which quantifies the efficiency of tidal energy dissipation in the star. Consequently, the orbital evolution timescale of hot Jupiters is strongly dependent on the adopted value of $Q^{\prime}_{\star}$. By substituting the appropriate $Q^{\prime}_{\star}$ value (see Section VI.3) along with other system parameters from Wong et al. (2022) into Equation (5) of Levrard et al. (2009):
@@
we estimate the remaining lifetime of WASP-12 b to be $\sim 0.41$ Myr, where $n=2\pi/P$ denotes the mean orbital motion of the planet.
VI.5 Estimation of Planetary Love Number
Tidal evolution theory predicts that hot Jupiters should circularize on timescales much shorter than the ages of their host stars (Levrard et al. 2007; Dawson and Johnson 2018). Assuming a planetary tidal dissipation factor of $Q_{p}\sim 10^{6}$ and applying the formalism of Patra et al. (2017), we estimate the tidal circularization timescale of WASP-12 b to be $\sim$ 0.34  Myr. This timescale is several orders of magnitude shorter than the estimated stellar age (WASP-12: $\sim$3.05  Gyr, Leonardi et al. 2024a), indicating that the planet’s orbit should have long been circularized. Any residual orbital eccentricity would therefore require additional mechanisms, such as ongoing perturbations or a re-evaluation of tidal dissipation efficiency, to be sustained (Bouma et al. 2019; Maciejewski et al. 2021).
A plausible explanation is that apsidal precession is influenced by the planet’s internal structure, as proposed by Ragozzine and Wolf (2009), since the precession rate depends on the tidal Love number ($k_{p}$), which encodes information about the planet’s internal density distribution. Following the formulation of Patra et al. (2017):
@@ VI.5 Estimation of Planetary Love Number
we estimated $k_{p}$ by substituting the values of $d\omega/dE$ (Table 8) together with other system parameters from Wong et al. (2022). This yields $k_{p}=0.63\pm 0.089$, a value consistent with Jupiter’s tidal Love number ($k_{p}=0.59$; Wahl et al. 2016). This similarity suggests that the interior density profile of WASP-12 b may be comparable to that of Jupiter. To test this hypothesis, continued monitoring of future transits and occultations will be essential.
VII Conclusions
@@
Previous studies of WASP-12 b (Maciejewski et al., 2016a; Patra et al., 2017; Yee et al., 2020) have revealed intriguing aspects of the system’s orbital dynamics. Early investigations reported the presence of short-term transit timing deviations potentially caused by an additional body (Maciejewski et al., 2011b), while later studies identified evidence of long-term timing deviations. Specifically, Maciejewski et al. (2016a); Patra et al. (2017); Maciejewski et al. (2018) observed a declining trend in transit timings indicative of orbital decay, whereas Bailey and Goodman (2019) proposed apsidal precession as a plausible mechanism. More recently, Yee et al. (2020) provided strong evidence supporting orbital decay in the WASP-12 b system. Detecting such orbital evolution requires long-term, high-precision monitoring. In this work, we combined newly acquired ground-based photometric data from 7 nights, obtained at various observatories worldwide, with publicly available light curves from the literature, ETD, ExoClock, and high-cadence, high-precision space-based observations from TESS to construct an extensive dataset for re-examining this hot Jupiter.
@@
We included a total of 119 complete transit light curves of WASP-12 b observed by TESS across six sectors (20, 43, 44, 45, 71, and 72). To extend the observational time baseline, we incorporated an additional 131 high-quality transit light curves from public databases, including ETD and ExoClock, along with 108 publicly available complete transit observations. Furthermore, 26 transit light curves were obtained via private communication with the respective authors and we also included 7 new ground based observations in this study. In total, 391 light curves were modeled to refine the system’s physical and orbital parameters and to determine individual mid-transit times. To ensure uniformity and precision, all light curves were analyzed using a consistent methodology for deriving the mid-transit times.
@@
Our transit-timing analysis indicates that the orbit of WASP-12 b is decaying at a rate of $31.97\pm 0.80~\mathrm{ms\,yr^{-1}}$ corresponding to a modified stellar tidal quality factor of $Q^{\prime}_{\star}=(1.52\pm 0.038)\times 10^{5}$. Statistical diagnostics, including reduced $\chi^{2}$, BIC, and AIC, favor the orbital decay model as the most plausible explanation for the observed timing variations.
@@
To constrain the planet’s internal structure, we calculated the planetary tidal Love number, obtaining $k_{p}=0.63\pm 0.089$. This value is consistent with Jupiter’s Love number ($k_{p}=0.59$; Wahl et al. 2016), suggesting a similar internal density distribution for WASP-12 b. Continued long-term, high-precision monitoring of future transits with forthcoming missions such as PLATO (Rauer et al. 2025) and ARIEL (Morales et al. 2022), complemented by secondary-eclipse (occultation) observations, will be critical for identifying the physical mechanism responsible for the observed timing offsets in WASP-12 b.
@@
We would like to thank the anonymous reviewer for his/her valuable suggestions and constructive feedback, which have substantially enhanced the quality of this paper. I.G.J. acknowledges funding from the National Science and Technology Council (NSTC), Taiwan, through Grants NSTC 113-2112-M-007-030, NSTC 114-2112-M-007-029, and NSTC 113-2115-M-007-008. This work utilizes new photometric observations obtained with the 0.61 m VASISTHA Telescope at the Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), Sitapur, Paschim Midnapore, operated by ICSP. We are grateful to Prof. Sandip Kumar Chakraborty for providing access to this facility and to the IERCOO staff members for their invaluable support during the observing runs. Additionally, this study incorporates observations from the 1.3 m Devasthal Fast Optical Telescope (DFOT) at the Devasthal Campus of ARIES, Nainital, and the 0.3 m AG Optical telescope at the Utah Desert Remote Observatory (UDRO) in Beryl, Utah. We thank the staff at DFOT and UDRO for their assistance. We also acknowledge Hebb Lesbie, Efrain Alavarado, and Karen Collins for sharing their transit light curves through private communication, as well as all other authors who made their published light curves publicly available, which proved to be an invaluable resource for this study. This work also makes use of data from the Transiting Exoplanet Survey Satellite (TESS) mission, publicly available through the Mikulski Archive for Space Telescopes (MAST). We gratefully acknowledge the TESS mission for its significant contribution to exoplanet science by providing high-precision transit photometric data. The specific observations analyzed in this study can be accessed via the TESS LCs–All Sectors page (MAST Team 2021 ), and funding for the mission is provided by NASA’s Explorer Program. Furthermore, this study utilized publicly available transit light curves from the Exoplanet Transit Database (ETD) and the ExoClock Project; we thank the contributors of these initiatives for making their data accessible. Currently, all the original lightcurves of ETD are available in VarAstro Server. Finally, this work made use of the NASA Exoplanet Archive, operated by the California Institute of Technology under contract with NASA as part of the Exoplanet Exploration Program.